OpenStudy (anonymous):

Each marble bag sold by Martina's Marble Company contains 4 green marbles for every 3 yellow marbles. If a bag has 24 green marbles, how many yellow marbles does it contain?

OpenStudy (anonymous):



OpenStudy (anonymous):

x = 18. so 18 yellow marbles (pretty sure)

OpenStudy (anonymous):

you can do it slowly with :4g+3y. =24. if we have one set:4+3=7. if we have 2 sets: 3*2 (6) yellows. 4*2(=8) 6+8=14. three sets...